OpenStudy (anonymous):

CHECK PLEASE! Assume the triangle has the given measurements. Solve for the remaining sides and angles. C = 95º , a = 1/2 , and b = 1/3

OpenStudy (anonymous):



OpenStudy (zzr0ck3r):

use law of cosines to get a \[\sqrt{(\frac{1}{2})^2+(\frac{1}{3})^2-2\frac{1}{2}\frac{1}{3}cos(95)}=a\]

OpenStudy (anonymous):

Okay. Yes that what I have! a = 0.6246

OpenStudy (zzr0ck3r):

now use law of sines to get A \[\frac{sin(95)}{0.6246}=\frac{sin(\theta)}{\frac{1}{2}}\]

OpenStudy (anonymous):

Oh okay nevermind then... And I do that for 1/3 also?

OpenStudy (anonymous):

Or should I just subtract my findings?

OpenStudy (zzr0ck3r):

either way

OpenStudy (zzr0ck3r):

those two equations wont help, try it and you will se what I mean

OpenStudy (anonymous):

Oh damn. Yes, I see now....

OpenStudy (anonymous):

This is what I have found so far 1.09392 = 2 sin(theta)

OpenStudy (anonymous):

And the other angle is 1.09392 = 3 sin(theta)

OpenStudy (anonymous):

@ankit042 @agent0smith @genius12 @Luigi0210 @Luis_Rivera How do I go on from here?

OpenStudy (agent0smith):

1.09392 = 2 sin(theta) If this is what you have, do you just need help solving for theta...?

OpenStudy (anonymous):

Yes! How do I find the angle for A and B?

OpenStudy (anonymous):

I know that 1.09392 = 2 sin(theta) is for A while 1.09392 = 3 sin(theta) is for B!

OpenStudy (agent0smith):

\[\Large \frac{\sin(95)}{0.6246}=\frac{\sin A }{\frac{1}{2}}\]use this for A.

OpenStudy (anonymous):

Oh wait, so thats it? Is there no way to break it down any further?

OpenStudy (agent0smith):

Put your calculator into degrees mode, it's in radians and your numbers aren't right.

OpenStudy (agent0smith):

1.09392 = 2 sin(theta) isn't correct... you used radians, the 95 is in degrees.

OpenStudy (anonymous):

Now I have 1.59493 = 3 sin(B) and 1.59493 = 2 sin(A)

OpenStudy (agent0smith):

Solving for A and B is easy now. For 1.59493 = 3 sin(B) , divide both sides by 3, then use inverse sine (sin^-1) 1.59493 = 2 sin(A) divide both sides by 2, then use inverse sine.

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

I got A = 52.888705 and B = 32.11655562

OpenStudy (agent0smith):

Those two add up to 85, and with the other angle being 95, the three add up to 180 degrees, which is a good sign.

OpenStudy (anonymous):

WOO-HOO!! Thank you!